Bethesda Mining Company reports the following balance sheet information for 2018 and 2019. BETHESDA MINING COMPANY Balance Sheets as of December 31, 2018 and 2019 2018 2019 2018 2019 Assets Liabilities and Owners’ Equity Current assets Current liabilities Cash $ 40,010 $ 51,181 Accounts payable $ 191,922 $ 199,611 Accounts receivable 60,281 80,639 Notes payable 87,020 138,588 Inventory 128,441 194,689 Total $ 278,942 $ 338,199 Total $ 228,732 $ 326,509 Long-term debt $ 241,000 $ 177,750 Owners’ equity Common stock and paid-in surplus $ 214,000 $ 214,000 Accumulated retained earnings 152,537 186,038 Fixed assets Net plant and equipment $ 657,747 $ 589,478 Total $ 366,537 $ 400,038 Total assets $ 886,479 $ 915,987 Total liabilities and owners’ equity $ 886,479 $ 915,987 Calculate the following financial ratios for each year: a. Current ratio. (Do not round intermediate calculations and round your answers to 2 decimal places, e.g., 32.16.) b. Quick ratio. (Do not round intermediate calculations and round your answers to 2 decimal places, e.g., 32.16.) c. Cash ratio. (Do not round intermediate calculations and round your answers to 2 decimal places, e.g., 32.16.) d. Debt-equity ratio and equity multiplier. (Do not round intermediate calculations and round your answers to 2 decimal places, e.g., 32.16.) e. Total debt ratio. (Do not round intermediate calculations and round your answers to 2 decimal places, e.g., 32.16.) Next Visit question mapQuestion 4 of 8 Total 4 of 8 Prev

Bethesda Mining Company reports the following balance sheet information for 2018 and 2019.
BETHESDA MINING COMPANY
Balance Sheets as of December 31, 2018 and 2019
2018 2019 2018 2019
Assets Liabilities and Owners’ Equity
  Current assets   Current liabilities
    Cash $ 40,010 $ 51,181      Accounts payable $ 191,922 $ 199,611
    Accounts receivable 60,281 80,639      Notes payable 87,020 138,588
    Inventory 128,441 194,689
            Total $ 278,942 $ 338,199
      Total $ 228,732 $ 326,509
  Long-term debt $ 241,000 $ 177,750
  Owners’ equity
     Common stock and paid-in surplus $ 214,000 $ 214,000
     Accumulated retained earnings 152,537 186,038
  Fixed assets   
    Net plant and equipment $ 657,747 $ 589,478            Total $ 366,537 $ 400,038
  Total assets $ 886,479 $ 915,987   Total liabilities and owners’ equity $ 886,479 $ 915,987
Calculate the following financial ratios for each year:
a. Current ratio. (Do not round intermediate calculations and round your answers to 2 decimal places, e.g., 32.16.)
b. Quick ratio. (Do not round intermediate calculations and round your answers to 2 decimal places, e.g., 32.16.)
c. Cash ratio. (Do not round intermediate calculations and round your answers to 2 decimal places, e.g., 32.16.)
d. Debt-equity ratio and equity multiplier. (Do not round intermediate calculations and round your answers to 2 decimal places, e.g., 32.16.)
e. Total debt ratio. (Do not round intermediate calculations and round your answers to 2 decimal places, e.g., 32.16.)

Next Visit question map

Question4of8Total4 of 8

Prev
Answer #1
2019 2018
Current ratio = current assets/current liabilities
=326509/338199 =228732/278942
                            0.97                            0.82
Quick Ratio = quick assets/current liabilities
=(current assets-inventories)/current liabilities
=(326509-194689)/338199 =(228732-128441)/278942
                            0.39                            0.36
Cash Ratio = Cash/current liabilities
=51181/338199 =60281/338199
                            0.15                            0.18
Debt equity ratio = total debt/total equity
=long term debt + current liabilities/total equity
=(177750+338199)/400038 =(241000+278942)/366537
                            1.29                            1.42
Equity multiplier = Total assets/Total equity
=915987/400038 =886479/366537
                            2.29                            2.42
Total debt ratio =total liabilities/total assets
=(long term debt+current liabilities)/total assets
=(177750+338199)/915987 =(241000+278942)/886479
                            0.56                            0.59